Use the simple interest formula to determine the missing value.
p=​$2,500, r=​?, t=88 ​years, i=​$880

Answer by fractalier(6550)

Here we can find r by dividing by Pt...so that
r = I/Pt = 880/(2500*88) = .004 = 0.4%